A resultant external force of 7 lbs acts on an object that
weighs 40 lbs on earth. What is the object’s acceleration on earth and on the
moon?
1
Expert's answer
2020-05-21T13:08:56-0400

The acceleration depends only on the mass of the object and the resultant external force.

Thus it will be equal in both cases.


"a=\frac{F_{net}}{m}=\frac{F_{net}}{W}g"

"a=\frac{F_{net}}{m}=\frac{7}{40}9.8=1.7\frac{m}{s^2}"